OpenStudy (anonymous):

I could really use a walk through of the method for solving this Approximate the critical number of f(x) = 18x cos (x) on the interval (0,pi) round to three decimal places.

OpenStudy (anonymous):

first we need the derivative, it is \[18(\cos(x)-x\sin(x))\] by the product rule

OpenStudy (anonymous):

oh good I managed to get that far lol but that was pretty much it.

OpenStudy (anonymous):

critical point will be the numbers in the interval \((0,\pi)\) where \[\cos(x)-x\sin(x)=0\]

OpenStudy (anonymous):

there is no easy way to solve this

OpenStudy (anonymous):

you can try \[\cos(x)=x\sin(x)\] or \[\frac{\cos(x)}{x}=\sin(x)\] but algebra will not get it for you, you need some sort of technology or newton-ralphson method. i would use technology

OpenStudy (anonymous):

the ugly decimals will show you that there is no snap way to do it

OpenStudy (anonymous):

we were working on newton's method I ended up with 1.57 as an approximation of the zero but I was not sure if that was right.

OpenStudy (anonymous):

i guess not since that is neither of the answers from wolf

OpenStudy (anonymous):

even newton's method will require technolgy because there is no way for you know know what say \[\sin(1)\] is so forget that mess and cheat

OpenStudy (anonymous):

lol ok thanks I did get .86 with my calculator initially.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

good, others are there as well

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

the others aren't in my interval.

OpenStudy (anonymous):

really interesting...

OpenStudy (anonymous):

yeah only one is in the inteval